# Enabling Efficient Optimal Control of Reconfigurable Battery Systems by Unifying System Modeling and Narrowing Search Space

### v1

@@ IV-B Identifying all feasible system configurations
1:Define five cell position flags: the $M_{p}$ at the beginning of a $C_{ps}$ ($FP$), the $M_{p}$ at the middle of a $C_{ps}$ ($MP$), the $M_{p}$ at the end of a $C_{ps}$ ($LP$), a $M_{s}$ ($S$), and bypassing ($B$). Represent a $C_{ps}$ by a vector consisting of all cell flags.
2:Initialize $D$ as a $N_{B}\times N_{B}\times 3$ list of cell flag vectors, and let $D[i][j][:]$ represent the flag vectors with a total of $i$ cells and meeting a terminal voltage level $j$.
3:If the last non-bypassed cell’s position flag in a flag vector is $FP$, $S$, $MP$, or $LP$, place the flag vector into the sub-list $D[i][j][0]$, $D[i][j][1]$, $D[i][j][1]$, or $D[i][j][2]$, respectively.
4:Initialize $D[1][1][0]\leftarrow[[FP]]$
5:Initialize $D[1][1][1]\leftarrow[[S]]$
6:Initialize $D[1][1][2]\leftarrow[]$
7:for $i\leftarrow 2$ to $N_{B}$ do
8:     for $j\leftarrow 1$ to $i$ do
9:         if $j=1$ then
@@
10:              Initialize $all\_bypass\leftarrow[B]\times(i-1)$
11:              $D[i][j][0].append(all\_bypass.append(FP))$
12:              $D[i][j][1].append(all\_bypass.append(S))$
13:         end if
14:          $D[i][j][0].extend([h.append(B)\mid h\in D[i-1][j][0]])$
15:          $D[i][j][0].extend([h.append(FP)\mid h\in D[i-1][j][0]])$
16:          $D[i][j][1].extend([h.append(B)\mid h\in D[i-1][j][1]])$
17:          $D[i][j][1].extend([h.append(S)\mid h\in D[i-1][j-1][0]\text{ if }h.count(FP)>1])$
18:          $D[i][j][1].extend([h.append(S)\mid h\in D[i-1][j-1][1]])$
19:         for $n\_mp\leftarrow 2$ to $i-1$ do
20:               $D[i][j][1].extend([h.append(MP)\times n\_mp\mid h\in D[i-n\_mp][j-1][0]\text{ if }h.count(FP)>1])$
21:               $D[i][j][1].extend([h.append(MP)\times n\_mp\mid h\in D[i-n\_mp][j-1][1]])$
22:         end for
23:          $D[i][j][2].extend([h.append(B)\mid h\in D[i-1][j][2]])$
24:          $D[i][j][2].extend([h.append(LP)\mid h\in D[i-1][j-1][0]\text{ if }h.count(FP)>1])$
25:          $D[i][j][2].extend([h.append(LP)\mid h\in D[i-1][j-1][1]])$
26:          $D[i][j][2].extend([h.append(LP)\mid h\in D[i-1][j][2]])$
27:     end for
28:end for
29:return $D[N_{B}][L_{v_{t}}^{min}:L_{v_{t}}^{max}][:]$
@@
Finally, according to the mapping relation established in Table III, all the identified feasible $C_{sp}$’s and $C_{ps}$’s are translated to feasible SSVs, which compose the feasible search space for system SSVs $\mathcal{B}_{f}^{N_{S}}$. Due to the fact that RBS design (a) encompasses other designs (b)–(f) shown in Fig. 1, their feasible search space can be constructed by excluding the SSVs that assign a state of 1 to switches that are absent in the designs.
@@ IV-C Quantifying the narrowing effect of the proposed feasible search space
To demonstrate the narrowing effect of the identified feasible search space $\mathcal{B}_{f}^{N_{S}}$ as compared with the complete search space $\mathcal{B}^{N_{S}}$, the numbers of feasible SSVs for various cell numbers and terminal voltage levels, denoted by $N_{f}{(N_{B},L_{v_{t}})}$’s, are summarized and compared with the number of all possible SSVs in Table IV. For an $N_{B}$-cell RBS, the total number of feasible SSVs across all terminal voltage levels is $\sum_{L_{v_{t}}=1}^{N_{B}}N_{f}{(N_{B},L_{v_{t}})}$. For the same system size $N_{B}$, the total number of possible SSVs is $2^{5N_{B}-3}$ since each cell is associated with 5 switches except that the last one has only 2 switches as shown in Fig. 1 (a). As can be seen in Table IV, both the total number of feasible SSVs in $\mathcal{B}_{f}^{N_{S}}$ and the total number of possible SSVs in $\mathcal{B}^{N_{S}}$ grows as the system size $N_{B}$ increases, but the former is much smaller than the latter and their ratios get even smaller as $N_{B}$ increases. Thus, by narrowing down the search space to only those feasible SSVs, the search becomes much more efficient as the system scales up and the ``````` `curse of dimensionality"""" can be significantly alleviated.
TABLE IV: Comparison of the number of feasible SSVs $N_{f}(N_{B},L_{v_{t}})$ for RBSs with various cell numbers and terminal voltage levels.
$N_{f}{(N_{B},L_{v_{t}})}$ $N_{B}=2$ $N_{B}=3$ $N_{B}=4$ $N_{B}=5$ $N_{B}=6$ $N_{B}=7$ $N_{B}=8$ $N_{B}=9$ $N_{B}=10$
$L_{v_{t}}=1$ 3 7 15 31 63 127 255 511 1023
$L_{v_{t}}=2$ 1 5 18 54 144 356 839 1919 4307
$L_{v_{t}}=3$ 1 7 30 103 310 853 2200 5410
$L_{v_{t}}=4$ 1 9 47 187 631 1907 5327
$L_{v_{t}}=5$ 1 11 68 312 1186 3959
$L_{v_{t}}=6$ 1 13 93 485 2063
$L_{v_{t}}=7$ 1 15 122 714
$L_{v_{t}}=8$ 1 17 155
$L_{v_{t}}=9$ 1 19
$L_{v_{t}}=10$ 1
No. of feasible SSVs = $\sum_{L_{v_{t}}=1}^{N_{B}}N_{f}{(N_{B},L_{v_{t}})}$ 4 13 41 125 369 1062 2999 8348 22978
@@
$\sum_{L_{v_{t}}=1}^{N_{B}}N_{f}{(N_{B},L_{v_{t}})}/2^{5N_{B}-3}$ $3.125\times 10^{-2}$ $3.174\times 10^{-3}$ $3.128\times 10^{-4}$ $2.980\times 10^{-5}$ $2.749\times 10^{-6}$ $2.473\times 10^{-7}$ $2.182\times 10^{-8}$ $1.898\times 10^{-9}$ $1.633\times 10^{-10}$
IV-D Case study
To evaluate the search efficiency within the orginal search space $\mathcal{B}^{N_{S}}$ and the narrowed search space $\mathcal{B}_{f}^{N_{S}}$, a 10-cell RBS with design (d) in Fig. 1 is taken as an example. The RBS is discharged with a constant power of 70 W for 20 steps, and each step lasts 20 seconds. The initial SoCs for the 10 cells have a total SoC imbalance $SoC_{ibl}^{tot}=0.30$, in which $SoC_{ibl}^{tot}$ is defined as follows.
$SoC_{ibl}^{tot}=\sum_{n=1}^{N_{B}}\text{{SoC}}_{n}-N_{B}\times\min_{n=1}^{N_{B}}\text{{SoC}}_{n}.$ (32)
According to (32), when the system operation must be terminated because the minimum cell SoC has reached its lower limit for all cells, $SoC_{ibl}^{tot}$ represents the total cell SoCs that remain unused.
Then, an RBS optimal control problem is formulated for minimizing the total SoC imbalance $SoC_{ibl}^{tot}$ at the end of the last step. To solve this problem, the Genetic Algorithm (GA) [23] is deployed, in which the population size, crossover probability, and mutation probability are set to 100, 0.8, and 0.1, respectively. The narrowed search space is examined in comparison with the complete search space composed of all possible SSVs. The evolutions of $SoC_{ibl}^{tot}$ over generations using these two types of search space are compared in Fig. 9.
@@ IV-D Case study
It can be seen from Fig. 9 that the obtained total SoC imbalance $SoC_{ibl}^{tot}$ using the narrowed search space converges quickly to 0.031 after 220 generations. However, within the same number of generations of search in the complete search space, the $SoC_{ibl}^{tot}$ remains unchanged. Since there exist massive infeasible SSVs in the complete space, leading to short circuits, open circuits, and unacceptable system terminal voltage levels, it becomes difficult to select feasible actions consecutively in the GA, forcing the search to end prematurely. This case study shows the proposed algorithm for narrowing the search space to only feasible solution set can facilitate the fast convergence to the optimal solution.
@@ V Conclusions
This study aims to address two fundamental challenges in the optimal control of RBSs: constructing the system plant expression during problem formulation and enhancing search efficiency during problem-solving.
The proposed unified modeling framework eliminates the need to frequently reconstruct the system plant during dynamic battery reconfiguration. This enables the formulation of optimal control problems for RBSs while significantly reducing computational time and storage requirements. Compared to experimental test results, the proposed RBS model demonstrates high accuracy in simulating dynamic current and voltage distributions within the system.
@@
Building on these essential techniques for RBSs, a variety of optimal control problems can be formulated and efficiently solved to enhance system performance from multiple perspectives, including improving fault tolerance, optimizing energy storage and delivery, enhancing charge balance, and extending the remaining useful life.

### v2

@@ IV-B Identifying all feasible system configurations
1:Define five cell position flags: in the $M_{p}$ at the beginning of a $C_{ps}$ ($P_{\text{bgn}}$), in any $M_{p}$ at the middle of a $C_{ps}$ ($P_{\text{mid}}$), in the $M_{p}$ at the end of a $C_{ps}$ ($P_{\text{end}}$), in any $M_{s}$ ($S$), and bypassed ($B$). Then, represent a $C_{ps}$ by a vector consisting of all cells’ flags, denoted by $V_{cp}$.
2:Initialize $VL$ as a $N_{B}\times N_{B}\times 6$ list of cell flag vectors, and let $VL[i][j][:]$ represent the list of $V_{cp}$’s with a total of $i$ cells and reaching the normalized system voltage $j$. If the last non-bypassed cell’s position flag in a $V_{cp}$ with $i$-cell and meeting $j$-level is $P_{\text{bgn}}$, $S$, or $P_{\text{end}}$, the $V_{cp}$ will be placed into the sub-list $VL[i][j][0]$, $VL[i][j][1]$, or $VL[i][j][2]$, respectively. The other sub-lists store transitional $V_{cp}$’s established only for augmentation. If all cells in the last module are labeled by $P_{\text{mid}}$, one single cell labeled by $P_{\text{bgn}}$ constitutes the first module, or one single cell labeled by $P_{\text{end}}$ constitutes the last module, the corresponding transitional $V_{cp}$ will be placed in $VL[i][j][3]$, $VL[i][j][4]$, or $VL[i][j][5]$, respectively.
3:Initialize $VL[1][1][1]\leftarrow[[S]]$
4:Initialize $VL[1][1][4]\leftarrow[[P_{\text{bgn}}]]$
5:for $i\leftarrow 2$ to $N_{B}$ do
6:     for $j\leftarrow 1$ to $i$ do
7:         if $j=1$ then
@@
8:              Initialize $all\_bypass\leftarrow[B]\times(i-1)$triangleright Note: $all\_bypass$ is a $V_{cp}$ with all $i-1$ entries equal to $B$.
9:              $VL[i][j][1].append(all\_bypass.append(S))$
10:              $VL[i][j][4].append(all\_bypass.append(P_{\text{bgn}}))$
11:         end if
12:          $VL[i][j][0].extend([h.append(B)\mid h\in VL[i-1][j][0]])$triangleright Note: $h$ represents any $V_{cp}$ extracted from a sub-list.
13:          $VL[i][j][0].extend([h.append(P_{\text{bgn}})\mid h\in VL[i-1][j][0]~{}\text{or}~{}VL[i-1][j][4]])$
14:          $VL[i][j][1].extend([h.append(B)\mid h\in VL[i-1][j][1]])$
15:          $VL[i][j][1].extend([h.append(S)\mid h\in VL[i-1][j-1][0],~{}VL[i-1][j-1][1],~{}\text{or}~{}VL[i-1][j-1][3]])$
16:          $VL[i][j][2].extend([h.append(B)\mid h\in VL[i-1][j][2]])$
17:          $VL[i][j][2].extend([h.append(P_{\text{end}})\mid h\in VL[i-1][j][2]~{}\text{or}~{}VL[i-1][j][5]])$
18:          $VL[i][j][3].extend([h.append(B)\mid h\in VL[i-1][j][3]])$
19:         for $n\_mp\leftarrow 2$ to $i-1$ do $\triangleright$ Note: $n\_mp$ is the number of cells in the $M_{p}$ at the middle of a $C_{ps}$.
20:               $VL[i][j][3].extend([h.append(P_{\text{mid}})\times n\_mp\mid h\in VL[i-n\_mp][j-1][0],~{}VL[i-n\_mp][j-1][1]~{}\text{or}~{}VL[i-n\_mp][j-1][3]])$
21:         end for
22:          $VL[i][j][4].extend([h.append(B)\mid h\in VL[i-1][j][4]])$
23:          $VL[i][j][5].extend([h.append(B)\mid h\in VL[i-1][j][5]])$
24:          $VL[i][j][5].extend([h.append(P_{\text{end}})\mid h\in VL[i-1][j-1][0],~{}VL[i-1][j-1][1]~{}\text{or}~{}VL[i-1][j-1][3]])$
25:     end for
26:end for
27:return $VL[N_{B}][\tilde{v}_{t}^{min}:\tilde{v}_{t}^{max}][0:2]$
As shown in Line 1 in Algorithm 1, the cell-module P&CPs listed in Table III are further aggregated, and five types of position flags of battery cells are defined, including the cell bypassing. Then, a $C_{ps}$ can be described by a vector consisting of all cells’ position flags, denoted by $V_{cp}$. To accommodate all $V_{cp}$’s, a three-dimensional list $VL$ is constructed in Line 2. The first dimension $i$ describes the number of cells in the system. The second dimension $j$ specifies the required system voltage level. Moreover, six types of sub-lists are constructed based on the third dimension. Among them, sub-lists $VL[i][j][0]$, $VL[i][j][1]$, and $VL[i][j][2]$ store $V_{cp}$’s corresponding to $C_{ps}$’s consisting of only one $M_{p}$, having an $M_{s}$ at the end, or having an $M_{p}$ at the end, respectively. These sub-lists are separated since they need to be constructed in different ways. Sub-lists $VL[i][j][3]$, $VL[i][j][4]$, and $VL[i][j][5]$ store $V_{cp}$’s only used for transition, and, hence, they are not returned eventually.
Next, from Line 3 to Line 30, DP is deployed to iteratively search for the $V_{cp}$’s corresponding to all feasible $C_{ps}$’s. For single-cell subsystems, sub-lists $VL[1][1][1]$ and $VL[1][1][4]$ are initialized in Lines 3 and 4, respectively, in which the inner brackets denote a vector and the outer brackets denote a vector list. For any multi-cell subsystem with $i>1$, sub-lists $VL[i][1][1]$ and $VL[i][1][4]$ need to be initialized in Lines 9 and 10, respectively, based on the $all\-bypass$ vector defined in Line 8. By appending a new item with value $x$ to the end of an array $y$ through the $append$ operation, an augmented array $y.append(x)$ can be obtained. For example, in Line 9 a cell position flag of $S$ is appended to the end of the vector $all\-bypass$ to get an augmented vector $all\-bypass.append(S)$, which is then appended to the end of the vector list $VL[i][j][1]$.
Then, for each normalized system voltage $j\in[1,i]$ of the $i$-cell subsystem, six types of sub-lists are constructed in Lines 12 to 24 through the $extend$ operation. Given two vector lists $x$ and $y$, $y.extend(x)$ represents an augmented vector list combining all vectors in both lists. In particular, the expression $h.append(P_{\text{mid}})\times n_{mp}$ in Lines 20 describes an augmented vector by appending $n_{mp}$ $P_{\text{mid}}$ cell position flags to the original vector $h$, indicating that a $M_{p}$ composed of $n_{mp}$ cells is connected in series with the original system with $i-n_{mp}$ cells. All possible ways need to be considered to construct the $V_{cp}$’s for each type of sub-list. For instance, a $V_{cp}$ for the sub-list $VL[i][j][1]$ can be constructed by appending a flag $B$ to the end of any $V_{cp}$ in sub-list $VL[i-1][j][1]$ or a flag $S$ to the end of any $V_{cp}$ in sub-lists $VL[i-1][j-1][0]$, $VL[i-1][j-1][1]$, or $VL[i-1][j-1][3]$, as shown in Lines 14 or 15. Through such iterative $append$ and $extend$ operations, the feasible $C_{ps}$’s for the investigated $N_{B}$-cell system can be gradually approached. Finally, the first three types of sub-lists $VL[N_{B}][\tilde{v}_{t}^{min}:\tilde{v}_{t}^{max}][0:2]$ for the $N_{B}$-cell system meeting the required normalized system voltages $\tilde{v}_{t}^{req}\in[\tilde{v}_{t}^{min},\tilde{v}_{t}^{max}]$ are returned for enumerating all feasible $C_{ps}$’s of the RBS.
@@
According to the mapping relation established in Table III, all the identified feasible $C_{sp}$’s and $C_{ps}$’s are translated to feasible SSVs, which compose the feasible search space for system SSVs $\mathcal{B}_{f}^{N_{S}}$. Furthermore, since RBS designs (b)–(f) can be realized by removing certain switch branches in RBS design (a) as shown in Fig. 1, their feasible search spaces can be obtained based on design (a)’s feasible search space by excluding those SSVs assigning state value 1 to the already removed switch branches.
@@ IV-C Quantifying the narrowing effect of the proposed feasible search space
For an $N_{B}$-cell RBS with design (a) as shown in Fig. 1 (a), the total number of switches is $N_{S}=5N_{B}-3$ since each cell is associated with five switches except the last one that has only two switches. Then, the total number of possible SSVs in the complete search space $\mathcal{B}^{N_{S}}$ is $2^{N_{S}}=2^{5N_{B}-3}$. When constructing the feasible search space of SSVs, denoted by $\mathcal{B}_{f}^{N_{S}}$, both the system size and the required normalized system voltage need to be considered. Thus, denote the total number of feasible SSVs meeting the normalized system voltage $\tilde{v}_{t}$ in $\mathcal{B}_{f}^{N_{S}}$ by $N_{f}{(N_{B},\tilde{v}_{t})}$. Then, the total number of feasible SSVs across all normalized system voltages is $\sum_{\tilde{v}_{t}=1}^{N_{B}}N_{f}{(N_{B},\tilde{v}_{t})}$.
To demonstrate the narrowing effect of the identified feasible search space, the number of feasible SSVs in the $\mathcal{B}_{f}^{N_{S}}$ is compared with the number of all possible SSVs in the $\mathcal{B}^{N_{S}}$ in Table IV. The sizes of $\mathcal{B}_{f}^{N_{S}}$ and $\mathcal{B}^{N_{S}}$, in terms of the number of SSVs, both grow as the system size $N_{B}$ increases, but the former is much smaller than the latter and their ratios get even smaller as $N_{B}$ increases. Taking the 10-cell system as an example, even though all normalized system voltages are allowed, the size ratio of $\mathcal{B}_{f}^{N_{S}}$ and $\mathcal{B}^{N_{S}}$ is very tiny, less than $2\times 10^{-10}$. Thus, the search effort is expected to be significantly reduced by narrowing the search space to only feasible SSVs, potentially enabling the solution of optimal control problems for RBSs. A case study is provided to demonstrate the effectiveness of the narrowed search space.
TABLE IV: Comparison of the number of feasible SSVs $N_{f}(N_{B},\tilde{v}_{t})$ for RBSs with various cell numbers and normalized system voltages.
$N_{f}{(N_{B},\tilde{v}_{t})}$ $N_{B}=2$ $N_{B}=3$ $N_{B}=4$ $N_{B}=5$ $N_{B}=6$ $N_{B}=7$ $N_{B}=8$ $N_{B}=9$ $N_{B}=10$
$\tilde{v}_{t}=1$ 3 7 15 31 63 127 255 511 1023
$\tilde{v}_{t}=2$ 1 5 18 54 144 356 839 1919 4307
$\tilde{v}_{t}=3$ 1 7 30 103 310 853 2200 5410
$\tilde{v}_{t}=4$ 1 9 47 187 631 1907 5327
$\tilde{v}_{t}=5$ 1 11 68 312 1186 3959
$\tilde{v}_{t}=6$ 1 13 93 485 2063
$\tilde{v}_{t}=7$ 1 15 122 714
$\tilde{v}_{t}=8$ 1 17 155
$\tilde{v}_{t}=9$ 1 19
$\tilde{v}_{t}=10$ 1
No. of feasible SSVs = $\sum_{\tilde{v}_{t}=1}^{N_{B}}N_{f}{(N_{B},\tilde{v}_{t})}$ 4 13 41 125 369 1062 2999 8348 22978
@@
$\sum_{\tilde{v}_{t}=1}^{N_{B}}N_{f}{(N_{B},\tilde{v}_{t})}/2^{5N_{B}-3}$ $3.125\times 10^{-2}$ $3.174\times 10^{-3}$ $3.128\times 10^{-4}$ $2.980\times 10^{-5}$ $2.749\times 10^{-6}$ $2.473\times 10^{-7}$ $2.182\times 10^{-8}$ $1.898\times 10^{-9}$ $1.633\times 10^{-10}$
IV-D Case study for comparing the searching effects within the complete and feasible search spaces
To compare the searching effects within the complete search space $\mathcal{B}^{N_{S}}$ and the narrowed search space $\mathcal{B}_{f}^{N_{S}}$, a 10-cell RBS with design (d) in Fig. 1 is taken as an example. The RBS is discharged with a constant power of 70 W for 20 decision steps, and each decision step lasts 20 seconds. For each battery cell, the SoC is constrained between 0.05 and 1, and the discharge current rate can reach up to 6 C (corresponding to 12 A). The normalized system voltage is restricted to at least four (i.e., at least four cells or modules need to be connected in series). The initial SoCs for the 10 cells have a total SoC imbalance $SoC_{ibl}^{tot}=0.30$, which is defined as follows.
$SoC_{ibl}^{tot}=\sum_{n=1}^{N_{B}}\text{{SoC}}_{n}-N_{B}\min_{n=1}^{N_{B}}\text{{SoC}}_{n}.$ (32)
In particular, $SoC_{ibl}^{tot}$ represents the total unused SoC when any cell SoC reaches the lower limit, necessitating the termination of system operation. Then, an RBS optimal control problem is formulated for minimizing the $SoC_{ibl}^{tot}$ at the end of the last decision step. A commonly used heuristic method, Genetic Algorithm (GA) [23], is deployed to solve this problem, in which the population size of individuals in each generation, the number of generations, crossover probability, and mutation probability are set to 100, 220, 0.8, and 0.1, respectively.
Both the complete search space and the feasible search space are tested when solving the problem, and the resulting evolutions of $SoC_{ibl}^{tot}$ over generations are compared in Fig. 9. When searching within the complete search space, the $SoC_{ibl}^{tot}$ remains unchanged over successive generations, indicating that the problem remains unsolved. The primary reason is that the majority of SSVs in the complete search space are infeasible for each decision step, as shown in Table IV. This makes it extremely challenging to select a feasible SSV for the initial decision step, let alone a series of consecutive feasible actions across all 20 decision steps. Indeed, it has been found that the SSVs generated for the first decision steps of all 100 individuals in each generation are infeasible. Consequently, for each generation, there is no need to evaluate the balancing performance of each individual across all decision steps, and the $SoC_{ibl}^{tot}$ is not improved over generations. This challenge can be addressed if the search space is narrowed to only feasible SSVs. As a result, the GA using the feasible search space manages to reduce the $SoC_{ibl}^{tot}$ from 0.3 to 0.031 after 220 generations. Thus, the proposed feasible search space enables the convergence to the optimal solution. Specifically, the optimized system configuration for each decision step is described in Fig. 10 by battery cells’ connection patterns within their associated modules, i.e., series, parallel, or bypassed. Moreover, the corresponding evolutions of cell SoCs over decision steps are presented in Fig. 11. Clearly, the cell SoC imbalance gradually decreases over decision steps due to the optimized system configurations illustrated in Fig. 10.
@@ IV-D Case study for comparing the searching effects within the complete and feasible search spaces
Figure 10: Optimized system configurations described by the cell connection patterns during all decision steps .
Figure 11: The cell SoC evolutions over decision steps under the optimized system configurations.
@@ V Conclusions
This study aims to address two fundamental challenges in the optimal control of RBSs: constructing the system model expression during problem formulation and enhancing search efficiency during problem-solving.
The proposed unified modeling framework based on the RBS’s CSC eliminates the need to frequently reconstruct the system model during dynamic battery reconfiguration. This enables the formulation of optimal control problems for RBSs while significantly reducing computational time and storage requirements. Compared to experimental test results, the proposed RBS model demonstrates high accuracy in simulating dynamic current and voltage distributions within the system.
@@
Building upon these essential techniques for RBSs, a variety of optimal control problems can be formulated and efficiently solved by selecting appropriate optimization algorithms to enhance system performance from multiple perspectives. This will be further investigated in our future study.